OpenStudy (anonymous):

if x^5 +x^2y +7y^2 = 15, find y' by implicit differentiation. y'=

OpenStudy (turingtest):

have you tried taking the derivative yet? If so what did you get?

OpenStudy (paxpolaris):

\[\large x^5 +x^2y +7y^2 = 15\ \ or\ \ \large x^5 +x^{2y} +7y^2 = 15\ ??\]

OpenStudy (anonymous):

the 1st one

OpenStudy (paxpolaris):

\[x^5 +x^2y +7y^2 = 15\] Differentiating both sides:\[\large 5x^4+2xy+x^2 \cdot y'+14y \cdot y'=0\]

OpenStudy (paxpolaris):

\[\large \implies\left( x^2+14y \right)y'=-\left( 5x^4+2xy \right)\] \[\Large \implies y'=-{5x^4+2xy \over x^2+14y}\]

OpenStudy (anonymous):

thankss